The accompanying table lists body temperatures from 68 different randomly selected subjects measured at two different times iBody Temperature (degrees F) at 8 AM I 97.8 Body Temperature (degrees F) at 12 AM 98.9 98.41 981 98.1 97.7 98.8 98.6 96.81 979 98.3 98.2 96.8 97.8 ysl 98.6| 983 98.3|| 97.5 29 7 | 97.4 9.5| 8.2 98.5 98.1 98. 98.9|| 98.1| 98.|| 98.9 1993 98.7 3 999 7.
Answer #1

Sample mean of the difference using excel function AVERAGE(), x̅d = -0.8279

Sample standard deviation of the difference using excel function STDEV.S(), sd = 0.8313

Sample size, n = 68

α = 0.05

a) Null and Alternative hypothesis:

Ho : µd = 0

H1 : µd ≠ 0

Test statistic:

t = (x̅d)/(sd/√n) = (-0.8279)/(0.8313/√68) = -8.21

df = n-1 = 67

p-value :

Two tailed p-value = T.DIST.2T(ABS(-8.2133), 67) = 0.0000

Conclusion:  

Since the p-value is less than the significance level, Reject Ho. There is sufficient evidence to warrant rejection of the claim of no difference between body temperature measured at 8 AM and at 12 AM.

b) 95% Confidence interval :

At α = 0.05 and df = n-1 = 67, two tailed critical value, t-crit = T.INV.2T(0.05, 67) = 1.996

Lower Bound = x̅d - t-crit*sd/√n = -0.8279 - 1.996 * 0.8313/√68 = -1.0291

Upper Bound = x̅d + t-crit*sd/√n = -0.8279 + 1.996 * 0.8313/√68 = -0.6267

-1.0291 < µd < -0.6267